how many liters each of 15% acid solution and 33% acid solution should be mixed to get 120 L of 21% acid solution?
what i have so far.
liter of sol. percent liter of acid
x 15% .15 .15x
y 33% .33 .33y
120 21% .21 .21(120)

Answer by stanbon(75887)

how many liters each of 15% acid solution and 33% acid solution should be mixed to get 120 L of 21% acid solution?
----
Equation:
acid + acid = acid
0.15x + 0.33(120-x) = 0.21*120
---
Multiply thru by 100 to get:
15x + 33*120 - 33x = 21*120
-----
-18x = -12*120
---
x = (2/3)*120
x = 80L (amt of 15% solution needed)
---
120-x = 40L (amt. of 33% solution needed)
============================================
Cheers,
Stan H.